What is the instantaneous velocity (in m/s) of a coconut falling vertically from a coconut tree after 2 2,0s?


1
Expert's answer
2021-07-27T16:40:01-0400

"u=0"

"t=22s"

"v=u+at"

"=0+(9.8\u00d722)=215.82m\/s"